• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 18:11:35 on September 29, 2010

    Entry number 333117

    keyword=Experiment status update (Wed)

    The CHL recovered fairly smoothly after yesterday's shutdown. The contamination has been removed and the Linacs are refilling. Beam should be available to the Halls by early swing.

    The ESR is also ready to go after a long down. That means the Techs can start cooling the LHRS, Moller, and Target. The LHRS will take 7--10 days to cool; the Moller should cool in 24 hours or so; the target will be cooled last. The basic plan for the next week will be to stay in Restricted Access during DAY shift so the Techs can push hard on cooling the Moller and LHRS. We will lock up during SWING and OWL and do what we can to commission the beamline, detectors, DAQ, and Moller. This includes BPM and BCM calibrations, beam optimization (eliminate beam halo), and anything else we can think of that does not require a functioning LHRS.

    For tonight (Wed), the rough plan is to lock up around 6--7pm and allow MCC to steer beam into the Hall through the Compton chicane down to the dump. Once that is complete and MCC hands over the beam to the Shift Leader, we can do a BPM calibration (ie. "Bull's Eye" scan), and take a quick look at backgrounds.

    I don't think there is any need to push things too late into the night. We should have plenty of time to get this kind of stuff done while the HRS cools.

    The Hall should return to Restricted Access at roughly 0700 tomorrow (Thursday) morning so the Techs can get back to cooling the magnets.

    Please maintain the Beam Time Accounting (BTA) numbers from now on. Make a note in the "Notes" field that the DAY shift downtime is associated with the LHRS cooling. I'm not sure how the adminstration will 'book' that downtime (against cryo/ESR I guess), but we should make a note of it on the BTA submissions.

    Here's a short list of suggested tasks for the week or so:

    A more extensive list is available here.

    A copy of this log entry has been emailed to: riordan, brads